A hollow rod of radius R1 and R2 such that R2 is greater than R1
Arnold solid rod of radius r are used to transmit power if they have the same mass and density which rod is stronger ?

Grade:11

2 Answers

Arun
4 years ago
Mass of the cylinder =  m
Outside radius = r₂
Inside radius = r₁
So the axis parallel to its symmetrical axis ( through its centre) and tangential to the outer surface means an axis on the outer surface through the length of the cylinder.
Moment of inertia on the symmetrical axis I₁ = (m/2) (r₁² + r₂²)
Following the parallel axis theorem, moment of inertia on the mentioned axis, I₂ = I₁ + mr₂² [ mass x distance from symmetrical axis ]
that is, I₂ = (m/2) (r₁² + r₂²) + mr₂²
or, I₂ = m [ (r₁²/2) + {(r₂²/2) + r₂²}
or I₂ = (m/2) [ r₁² + 3r₂² ]
So the answer is (m/2) (r₁² + r₂²)
Vikas TU
4 years ago
Dear student 
The ans is not complete
Arnold solid rod will be stronger. 
Hope this will help 
Good Luck